A hydraulic lift is designed to raise a 900 kg car
Answer #1

The weight of car = 900 kg * g = 900* 9.8 = 8820 N

To lift the car P1 = P2

area of piston 1 = pi*r12 = 3.14*0.352 = 0.38465 m2

area of piston 2 = pi*r22 = 3.14*0.022 = 0.001256 m2

force = preassure*area

which area has more we have to use that piston

i.e. 35 cm piston